CTO Transport has an aftertax cost of debt of 5.6 percent, a cost of equity of 13.7 percent, and a cost of preferred stock of 7.8 percent. The firm has 60,000 shares of common stock outstanding at a market price of $45 a share. There are 12,000 shares of preferred stock outstanding at a market price of $52 a share. The bond issue has a total face value of $400,000 and sells at 102 percent of face value. The tax rate is 35 percent. What is the company’s WACC?
Answer #1

Total value of common stock=(60000*45)=$2,700,000

Total value of preferred stock=(12000*52)=$624,000

Total value of bonds=400,000*102%=$408,000

Total value=(2,700,000+624000+408000)=$3,732,000

WACC=Respective costs*Respective investment weight

=(2,700,000/3,732,000*13.7)+(624000/3,732,000*7.8)+(408000/3,732,000*5.6)

=11.83%(Approx).